An artist wants her painting to be in the shape of a golden rectangle. If the length of the painting is 36 inches, then what should be the width?
Answer by ewatrrr(24785)


Hi
Let x represent the length of the width
A golden rectangle is one whose side lengths are in the golden ratio, or approximately 1:1.618.
1/1.618 = x/36in
1.618x = 36in
x = 36in/1.618

or 1/1.618 = 36in/x
x = 1.618*36in